OpenStudy (templeguy):

3. You are buying potato chips at the grocery store. You have a manufacturer’s coupon for $0.50 off and the store is offering 50% off if you buy 2 bags. How much is the total for 3 bags if each bag cost $2.50

OpenStudy (templeguy):

@quickstudent @whpalmer4 @retirEEd @gnech @HanaMalik @brainzonly

OpenStudy (danjs):

how many bags did you buy?

OpenStudy (templeguy):

I edited it

OpenStudy (danjs):

So you only get 50% off for multiples of 2 bags, the third one is full price still Total Cost = 2 bags * $2.50 * 0.50 + 1 bag * $2.50 - $0.50 coupon

OpenStudy (danjs):

Overall you have, 2 bags at half price, 1 full price bag, and the coupon

OpenStudy (templeguy):

7.75?

OpenStudy (templeguy):

@DanJS whats the answer?

OpenStudy (danjs):

That is it, just have to calculate it Total Cost = 2 bags * $2.50 * 0.50 + 1 bag * $2.50 - $0.50 coupon = 2*2.5*0.5 + 2.50 - 0.50 = 2.5 + 2.5 - 0.50 = 4.5